OpenStudy (steve816):

Please help me! The volume of a block of wood equals 78 cm^3 What is its volume in cubic inches?

OpenStudy (anonymous):

2.54cm / inch 78cm^3*(1 inch/2.54cm)^3 The cm cancel out =78/(2.54^3)

OpenStudy (steve816):

I don't think that is right...

OpenStudy (anonymous):

Stoichiometry has a fancy name but it's very easy. Just like \[\frac{ 1 }{ 2 }*\frac{3}{3} = \frac{3}{6} \] You are multiplying your value by 1, because the value of numerator equals the denominator. And just keep doing it so units cancel out to the ones you want to result to. 2.54 cm = 1 inch and you keep canceling the cm until you get in^3 (and not 1/in^3)

OpenStudy (anonymous):

\[78cm^3\frac{ 1 inch }{ 2.54cm }\frac{1 inch}{2.54cm}\frac{1inch}{2.54cm}\]

OpenStudy (anonymous):

It's correct. If you have cm in the numerator and multiply it by something that has cm in the denominator, the cms cancel out.

OpenStudy (anonymous):

Thanks kestry, it's tough entering fractions on a phone

OpenStudy (steve816):

Thanks everyone for the help :) I think I understand it now!

OpenStudy (anonymous):

Haha, you answered the question so fast and well!

OpenStudy (anonymous):

Woo, thanks steve! Good luck on the rest of your hw

OpenStudy (steve816):

Thanks kestry! By the way, what are you working on kestry?

OpenStudy (anonymous):

I'm working on Calculus /cry